How do I solve the problem:A box of goldfish food can feed 3 fish for 4 weeks. How long will the box last if there are 7 goldfish?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

A box of goldfish food can feed 3 fish for 4 weeks. How long will the box last if there are 7 goldfish?
-----
time is directly related to # of boxes of food and indirectly related to # of fish
--------------------------
Equation:
t = k*b/f
----
Solve for "k" using "A box of goldfish food can feed 3 fish for 4 weeks".
4 = k*1/3
---
k = 12
------
Equation:
t = 12b/f
---
How long will the box last if there are 7 goldfish?
t = 12b/f
t = 12*1/7
time = 1.71 weeks
=====================
Cheers,
Stan H.
=====================

Answer by richwmiller(17219)

3*7*4=84 portions
3*7*4/7=12 days for 7 fish